OpenStudy (anonymous):

4. Simplify the expression. -3(b – 7) This is what i got -3 -3b -21 +3 +3 -3 b -18 18/3 =6

OpenStudy (abdullahm):

Lets start from the beginning.

OpenStudy (abdullahm):

We need to use this property, which by chance I had told you in the earlier post. \(\sf\Large a(b+c)=ab+ac\)

OpenStudy (abdullahm):

HOWEVER, your question is \(\sf\Large -3(b-7)\)

OpenStudy (anonymous):

Yeah

OpenStudy (abdullahm):

so since you have a problem with negative numbers, let me make it a little easier on you :)

OpenStudy (abdullahm):

\(\sf\Large -a(b-c) = -ab + ac\) Remember to carefully look at the - and +'s :)

OpenStudy (anonymous):

Okay. Still don't really get it.

OpenStudy (abdullahm):

just replace the a with 3 and the c with 7

OpenStudy (anonymous):

-3(7) = -21?

OpenStudy (anonymous):

-3(-7) 21

OpenStudy (anonymous):

Is this right?

OpenStudy (abdullahm):

\(\sf\Large-3(b – 7)=\\\sf\Large (-3)(b) - (-3)(7)=? \)

OpenStudy (abdullahm):

and that is correct :)

OpenStudy (anonymous):

So i do need to do anything else with this problem?

OpenStudy (abdullahm):

but 21 isn't the final answer.

OpenStudy (abdullahm):

so far you have gotten: \(\sf\Large (-3)(b) - (-3)(7)=-3b-(-21)\)

OpenStudy (anonymous):

Why is it saying i'm typing and half the time i'm not. XD

OpenStudy (anonymous):

So do i solve for b?

OpenStudy (abdullahm):

that's an OpenStudy glitch xD

OpenStudy (abdullahm):

and no, just simplify that what is - (-21) = ?

OpenStudy (anonymous):

Positive 21?

OpenStudy (abdullahm):

correct

OpenStudy (abdullahm):

so your final answer would become: \(\sf\Large -3b+21\)

OpenStudy (anonymous):

Alright thank you so much ^ ^

OpenStudy (abdullahm):

:)

OpenStudy (anonymous):

Openstudy not letting me close the question

OpenStudy (abdullahm):

refresh, maybe?